OpenStudy (anonymous):

Solve (: 0.25(7 + 3m) = -0.125 m. A.2 B.-2 C.-40.5 D.40.5 Note-I can't reply

OpenStudy (anonymous):

You have to find the value of m, it's a linear equation (First order) 0.25=(1/4) 0.125=(1/8) \[\frac {1}{4} (7+3m)=- \frac {1}{8} m\] Multiply by 8 \[8*\frac {1}{4} (7+3m)=-8* \frac {1}{8} m\] \[2*(7+3m)=- m\] expand \[14+6m=-m\] rearrange 7m=-14 m=-2